# Asymmetric Errors

### v2

@@ 3.3 Goodness of fit with likelihood
[…]
Wilks’ theorem holds only in the limit of large $N$and provided the first model contains the second. The first point depends clearly on the number and precision of theresults being combined. . So this needs to be investigated empirically.
@@
It therefore appears that $\Delta\ln L$ – which emerges as part of the combination procedure – can be used as a criterion for the acceptability of the combination, though it should not be treated as an exact quantity as the conditions for the validity of Wilks’ theorem may not be met.
@@ 3.4 Combination of errors with likelihood
For the linear sigma model, $w_{i}={\left(\sigma_{i}+\sigma^{\prime}_{i}a_{i}\right)^{3}\over\sigma_{i}}$, for the linear variance model $w_{i}={\left(V_{i}+a_{i}V^{\prime}_{i}\right)^{2}\over 2V_{i}+V^{\prime}_{i}a_{i}}$, and similarly for other models. If the cancellation cannot be done algebraically and numerical evaluation gives 0/0 at the peak, one can use $w=-\left({\partial^{2}L\over\partial a^{2}}\right)^{-1}$ for small $a$.
@@ 4.1.2 A lifetime measurement: combination of results using likelihoods
Suppose an experiment measures a lifetime $\tau$ from 3 decays. The values happen to be 1.241, 0.592, and 0.988, in some time units. (These numbers were generated from an exponential distribution, ensuring that this example is realistic.) Maximising $\ln L$, which is $\ln{\rm exp}(-t/\tau)/\tau$, and using $\Delta\ln L=-{1\over 2}$ gives a result $\tau={0.940}^{+0.841}_{-0.385}$.
@@
If we combine all 6 values we get the combinedresult ${1.1325}^{+0.6225}_{-0.3598}$. The likelihood used is the product of 6 separate exponential likelihoods, so we are using the full information..
@@ 4.1.3 Combining Poisson counts: combination of results using likelihoods
${9}^{+3.342}_{-2.676}$ ${1}^{+1.358}_{-0.6983}$ ${7.348}^{+3.149}_{-2.549}$ ${5.201}^{+1.942}_{-1.605}$ ${3.633}^{+1.426}_{-1.414}$ ${3.209}^{+0.5942}_{-0.8277}$
Table 10: Combining results from two samples from the same Poisson distribution. The ideal result is ${5.000}^{+1.752}_{-1.419}$.
@@ 4.1.4 Strength of a known mass peak: models of a known likelihood
A common analysis method is to take measurements of a reconstructed particle mass, and fit them to a total of signal and background shapes where all parameters are known except for the normalisations. Figure 16 shows such a situation, where the background is known to be flat and the signal is a Gaussian of mean 5.0 and standard deviation 1.0. One may then, as appropriate, work with the raw (unbinned) likelihood $\ln L=\sum_{i}\ln{(N_{B}B(x_{i})+N_{S}S(x_{i}))}$, where the sum is over all events, or the binned likelihood $\ln L=\sum_{j}(F_{j}\ln n_{j}-F_{j})$where $n_{j}$ is the contents of bin $j$ and $F_{j}$ is the prediction. The likelihood curves are shown in the second and third panels of Figure 16. In this example the unbinned likelihood gives a result $N_{S}={39.41}^{+10.20}_{-10.13}$, and the binned likelihood gives $N_{S}={40.43}^{+10.40}_{-10.35}$.
@@
Figure 17: Modelling the likelihoods in the case shown in Figure 16, plotting the differences from the true curve.
@@ 4.1.5 A product with asymmetric uncertainties
The central value is $1000\times 12.3\times 0.12=1476$ events. Writing $\sigma=\sigma_{0}+\Delta_{\sigma},F=F_{0}+\Delta_{F}$ then, to leading order in a Taylor expansion, $N=N_{0}+({\cal L}F_{0})\Delta_{\sigma}+({\cal L}\sigma_{0})\Delta_{F}$, and the appropriate function (pdf or $\ln L$) is found by combining the errors, scaled by ${\cal L}F_{0}$ and ${\cal L}\sigma_{0}$ respectively. For the linear sigma model this gives ${1476}^{+136}_{-250}$. Using the linear variance model it is ${1476}^{+137}_{-251}$.
@@ 5.1 Framework
As mentioned in Section 1.2, a pdf can give 68% central limits, but they relate tothe horizontal lines on the Neyman confidence belt, not the vertical ones, and the relationship is subtle. In the example discussed in Figure 2 a pdf with positive skewness leads to a likelihood with negative asymmetry.Essentially the same problem occurs in the evaluation of confidence regions using the bootstrap. A pdf is approximatedby taking repeated samples from the data, and these can be used to give a confidence region for some quantity. However this is not the confidence region for the true quantity, and to use it as such “amounts to looking up the wrong tables backwards”, as Peter Hall puts it [19, p36] and [20, p928]. The proportional Gaussian: $p(x)={1\over\sqrt{2\pi}\alpha a}e^{-{1\over 2}\left({x-a\over\alpha a}\right)^{2}}$ has a symmetric pdf but a likelihood with positive asymmetry.The Poisson has a positive asymmetry for both probability and likelihood, as its variance increases with the mean. So there is no easy equivalence between the 68% confidence region obtained from the pdf and that from the likelihood.
5.2 What happens if you use the wrong formalism.
@@
As an example we consider the combination of errors from two results, both with error ${\sigma^{+}}=2,{\sigma^{-}}=1$. These are shown in Table 11 for several models. Not surprisingly – this is a large asymmetry – the different models give a spread of different values for the combined ${\sigma^{+}}$, and likewise for ${\sigma^{-}}$. But the pdf models are clustered separately from the $\ln L$ models. The values of ${\sigma^{+}}$ and ${\sigma^{-}}$ for the combination under the assumption that these come from pdfs are both larger than those under the assumption that these come from likelihoods. This discrepancy is not enormous, but it is larger than the differences between individual models. (The difference between any of these and the ‘wrong’ method of summing separately in quadrature is somewhat larger.) The conclusion to be tentatively drawn is that if a pdf error is (wrongly) treated as a likelihood error, or vice versa, this makes a meaningful, though not enormous, difference.
@@ 5.3 A final result
We suppose that $\eta$ and $S$ are well known, but that there is a Gaussian error $\sigma_{t}=0.1s$ on the calibration of the time measurement $t$ which is systematic, i.e. described by a pdf, and gives an asymmetric error on $1/t$.
Considering an example where $A=100,\eta=1,n=50$ and $t=1s$, we would write $R={5000}^{+741}_{-674}{}^{+556}_{-455}$ where the first, ’statistical’, error comes from the Poisson errors on $n=50$ and the second ’systematic’ from the error on $1/t$.
If we, wrongly, treat the pdf error as if it were a likelihood error and combine the two errors this gives a combined result of ${5000}^{+906}_{-828}$ (using the linear-sigma model: other models give near-identical values). If we, wrongly, treat the lnL error as if it were a pdf and combine them, this gives ${5000}^{+909}_{-831}$ (using the dimidiated model: other models give near-identical values. There is little difference between the two outcomes.
As a validation we performed 10,000,000 toy experiments, generating values of $n$ from a Poisson distribution of mean 50 and values of $t$ from a Gaussian distribution of mean 1 and standard deviation 0.1, and applying the two prescriptions. The pdf method gave a coverage of $69.074\pm 0.015$%, only slightly above the nominal 68.269%, and the lnL method was even closer with a coverage of $68.387\pm{0.015}$%
@@ 7 Acknowledgements
We would like to thank our colleagues for many discussions and suggestions, (including a PHYSTAT informal review), especially the PHYSTAT organisers: Lydia Brenner, Nick Wardle, Olaf Behnke, Sara Algeri and Louis Lyons (who suggested some of the examples). This paper arose out of a workshop at the Banff Center for Arts and Creativity in Alberta, Canada, and we are very happy to thank the Banff International Research Station (BIRS) and itsstaff for their organisation and hospitality. We are grateful to Hans Dembinski for helpful comments.
ARB acknowledges that the publication was produced with funding from the Italian Ministry of University and Research under the Call for Proposals related to the scrolling of the final rankings of the PRIN 2022 call – Project title “Latent variable models and dimensionality reduction methods for complex data” – Project No. 20224CRB9E, CUP C53C24000730006 – PI Prof. Paolo Giordani – Grant Assignment Decree No. 1401 adopted on 18.9.2024 by MUR.
@@ A.1 The dimidiated Gaussian
Although the model is clearly unrealistic, it has some nice simple features. The quantile parameters provide the parameters of the pdf directly, as given by Equation (16).
@@
To determine the quantile parameters from the moments these equations have to be solved numerically. If we write $D={\sigma^{+}}-{\sigma^{-}}$ and $S={\sigma^{+}}^{2}+{\sigma^{-}}^{2}$ the equations
@@
${|\gamma|\over\sqrt{V^{3}}}={\pi+1\over\sqrt{(\pi-1)^{3}}}\approx 1.641$ (20)
so asymmetries, in the $\gamma\over{V^{3/2}}$ sense larger than this cannot be handled. But such cases are extreme.
@@ A.2 The distorted Gaussian
A more sophisticated procedure to handle the results of an OPAT analysis is to draw a parabola through the three points, as in Figure 20.
@@
The parameters $a$ and $b$ are given very simply from the quantile parameters: $a={1\over 2}({\sigma^{+}}+{\sigma^{-}})$ and $b={1\over 2}({\sigma^{+}}-{\sigma^{-}})$. The pdf is again the probability transform of the unit Gaussian, but using a single quadratic transform rather than two piecewise linear ones. Although this is more plausible than the two straight lines of the dimidiated Gaussian, care needs to be taken to include both arms of the parabola in evaluating the pdf. The support for $p(R)$ is limited by the minimum (or, for a negative skewness, the maximum) of the parabola. Where this occurs there is a Jacobian peak, and this appears in Figure 20 around $R=22$.
If $\nu$ is distributed according to the unit Gaussian with $\mu=0$ and $\sigma=1$ and $x(\nu)=x_{0}+a\nu+b\nu^{2}$, then the pdf for $x$ is $p(x)={{\phi}(\nu)\over|x^{\prime}(\nu)|}={{\phi}({\nu})\over|a+2b\nu|}$, where $a$ and $b$ are as given in Equation (24), and $\phi(\nu)$ is the unit Gaussian. As this is a quadratic, there are two values of $\nu$ for a given $x$: $\nu={\sqrt{a^{2}+4bx}-a\over 2b}$ and ${-\sqrt{a^{2}+4bx}-a\over 2b}$. Except for small asymmetries and/or small deviations, where the contribution from the second can be neglected, the pdf for $x$ is the sum of the two.
@@ A.3 The “railway” Gaussian
The railway Gaussian density attempts to mitigate the drawbacks of the dimidiated Gaussian (the discontinuity in the middle) and of the distorted Gaussian (the Jacobian spike). This density is also based on a coordinate transformation. The curve drawn through the OPAT analysis points is a parabola in the $[-1,1]$ region which smoothly transitions into straight lines on both left and right sides of the region. These transition regions are highly reminiscent of the track transition curves in railroad tracks, and this similarity gives the density its name. The coordinate transformation is defined by the equation
@@
The R code (Appendix D) evaluates the moments of the railway Gaussian algebraically, whereas in the python/C++ code (Appendix C) the moments are evaluated numerically and, for numerical reasons, the values of $h$ used are restricted to the range $[0.1,10.0]$.
@@ A.4 The double cubic Gaussian
Figure 22: The Double Cubic Gaussian: the panel on the left illustrates the coordinate transformation used to construct a double cubic Gaussian with parameters $M=0$, $\sigma^{+}=1$, $\sigma^{-}=0.6$. The panel on the right shows the corresponding density.
@@ A.5 The symmetric beta Gaussian
Figure 23: The Symmetric Beta Gaussian: the panel on the left illustrates the coordinate transformations used to construct densities shown on the right. In all cases, $M=0$, $\sigma^{+}=1$, $\sigma^{-}=0.6$.
@@ A.6 The quantile variable width Gaussian
An example QVW Gaussian density and the corresponding transformation of the standard normal are shown in Figure 24
Figure 24: The QVW Gaussian: the panel on the right illustrates the density which corresponds to the asymmetric uncertainty result $0^{+1.0}_{-0.6}$. The panel on the left shows the corresponding coordinate transformation given by $x=q(\Phi(z))$.
@@ A.7 The Fechner distribution
Figure 25: The Fechner distribution: the panel on the right illustrates the probability density function which corresponds to the asymmetric uncertainty result $0^{+1.0}_{-0.7}$. (The result $0^{+1.0}_{-0.6}$ for which a number of other distributions is plotted in this appendix can not be represented by Fechner.) The panel on the left shows the corresponding coordinate transformation given by $x=q(\Phi(z))$.

### v3

@@ 3.3 Goodness of fit with likelihood
According to Wilks’ theorem, under the null hypothesis the improvement in likelihood due to the introduction of $n$ extra parameters is such that $-2\Delta\ln L$ is distributed according to a $\chi^{2}$ distribution. On that basis, the change in the log likelihood when using only 1 free parameter (the combined result) rather then $n$ fitted values (the individual results) gives a measure of goodness of fit such that $-2\Delta\ln L$ is $\chi^{2}$ with $n-1$ degrees of freedom.
Wilks’ theorem holds only in the asymptotic limit,and provided the first model is contained bythe second. The first point depends clearly on the number and precision of the results being combined. So this needs to be investigated empirically.
@@
It therefore appears that $\Delta\ln L$ — which emerges as part of the combination procedure — can be used as a criterion for the acceptability of the combination, though it should not be treated as an exact quantity as the conditions for the validity of Wilks’ theorem may not be met.
@@ 3.4 Combination of errors with likelihood
For the linear sigma model, $w_{i}={\left(\sigma_{i}+\sigma^{\prime}_{i}a_{i}\right)^{3}\over\sigma_{i}}$, for the linear variance model $w_{i}={\left(V_{i}+a_{i}V^{\prime}_{i}\right)^{2}\over 2V_{i}+V^{\prime}_{i}a_{i}}$, and similarly for other models. If the cancellation of the $a_{i}$ factorcannot be done algebraically and numerical evaluation gives 0/0 at the peak, one can use $w=-\left({\partial^{2}L\over\partial a^{2}}\right)^{-1}$ for small $a$.
@@ 4.1.2 A lifetime measurement: combination of results using likelihoods
Suppose an experiment measures a lifetime $\tau$ from 3 decays. The values happen to be 1.241, 0.592, and 0.988, in some time units. (These numbers were generated from an exponential distribution with lifetime 1.0, ensuring that this example is realistic.) Maximising $\ln L$, which is $\ln\prod{\rm exp}(-t_{i}/\tau)/\tau$, and using $\Delta\ln L=-{1\over 2}$ gives a result $\tau={0.940}^{+0.841}_{-0.385}$.
@@
If we combine all 6 values we get the combined result ${1.1325}^{+0.6225}_{-0.3598}$. The likelihood used is the product of 6 separate exponential likelihoods, so we are using the full information.
@@ 4.1.3 Combining Poisson counts: combination of results using likelihoods
${9}^{+3.342}_{-2.676}$ ${1}^{+1.358}_{-0.698}$ ${7.348}^{+3.149}_{-2.549}$ ${5.201}^{+1.942}_{-1.605}$ ${3.633}^{+1.426}_{-1.414}$ ${3.209}^{+0.594}_{-0.828}$
Table 10: Combining results from two samples from the same Poisson distribution. The exact result is ${5.000}^{+1.752}_{-1.419}$.
@@ 4.1.4 Strength of a known mass peak: models of a known likelihood
A common analysis method is to take measurements of a reconstructed particle mass, and fit them to a total of signal and background shapes where all parameters are known except for the normalisations. Figure 16 shows such a situation, where the background is known to be flat and the signal is a Gaussian of mean 5.0 and standard deviation 1.0. One may then, as appropriate, work with the raw (unbinned) likelihood $\ln L=\sum_{i}\ln{(N_{B}B(x_{i})+N_{S}S(x_{i}))}$, where the sum is over all events, or the binned likelihood $\ln L=\sum_{j}(F_{j}\ln n_{j}-F_{j})$, where $n_{j}$ is the contents of bin $j$ and $F_{j}$ is the prediction. The likelihood curves are shown in the second and third panels of Figure 16. In this example the unbinned likelihood gives a result $N_{S}={39.41}^{+10.20}_{-10.13}$, and the binned likelihood gives $N_{S}={40.43}^{+10.40}_{-10.35}$.
@@
Figure 17: Modelling the likelihoods in the case shown in Figure 16, plotting the differences from the true curve. The different models give very similar functional forms, especially in the central region, making the point that results will be robust under choice of model, at least in this example. If the actual forms are shown, rather than the differences, the distinctions are even less apparent.
@@ 4.1.5 A product with asymmetric uncertainties
The central value is 1000×12.3×0.12=14761000$\times$ 12.3$\times$ $0$ 12.3 × 0.12 = 1476 events. Writing $\sigma=\sigma_{0}+\Delta_{\sigma},F=F_{0}+\Delta_{F}$ then, to leading order in a Taylor expansion, $N=N_{0}+({\cal L}F_{0})\Delta_{\sigma}+({\cal L}\sigma_{0})\Delta_{F}$, and the appropriate function (pdf or $\ln L$) is found by combining the errors, scaled by ${\cal L}F_{0}$ and ${\cal L}\sigma_{0}$ respectively. For the linear sigma model this gives ${1476}^{+136}_{-250}$. Using the linear variance model it is ${1476}^{+137}_{-251}$.
@@ 5.1 Framework
As mentioned in Section 1.2, a pdf can give 68% central limits, but they relate to the horizontal lines on the Neyman confidence belt, not the vertical ones, and the relationship is subtle. In the example discussed in Figure 2 a pdf with positive skewness leads to a likelihood with negative asymmetry. Essentially the same problem occurs in the evaluation of confidence regions using the bootstrap. A pdf is approximated by taking repeated samples from the data, and these can be used to give a confidence region for some quantity. However this is not the confidence region for the true quantity, and to use it as such “amounts to looking up the wrong tables backwards”, as Peter Hall puts it [19, p36] and [20, p928]. The proportional Gaussian has a symmetric pdf, $p(x)={1\over\sqrt{2\pi}\alpha a}e^{-{1\over 2}\left({x-a\over\alpha a}\right)^{2}}$, but a likelihood with positive asymmetry. The Poisson has a positive asymmetry for both probability and likelihood, as its variance increases with the mean. So there is no easy equivalence between the 68% confidence region obtained from the pdf and that from the likelihood.
5.2 What happens if you use the wrong formalism
@@
As an example we consider the combination of errors from two results, both with error ${\sigma^{+}}=2,{\sigma^{-}}=1$. These are shown in Table 11 for several models. Not surprisingly — this is a large asymmetry — the different models give a spread of different values for the combined ${\sigma^{+}}$, and likewise for ${\sigma^{-}}$. But the pdf models are clustered separately from the $\ln L$ models. The values of ${\sigma^{+}}$ and ${\sigma^{-}}$ for the combination under the assumption that these come from pdfs are both larger than those under the assumption that these come from likelihoods. This discrepancy is not enormous, but it is larger than the differences between individual models. (The difference between any of these and the ‘wrong’ method of summing separately in quadrature is somewhat larger.) The conclusion to be tentatively drawn in this instance is that if a pdf error is (wrongly) treated as a likelihood error, or vice versa, this makes a meaningful, though not enormous, difference. But this should not be taken as being true in general, and individual cases will need to be justified.
@@ 5.3 A final result
[…]
@@ 7 Acknowledgements
We would like to thank our colleagues for many discussions and suggestions, (including a PHYSTAT informal review), especially the PHYSTAT organisers: Sara Algeri, Olaf Behnke, Lydia Brenner, Louis Lyons (who suggested some of the examples) and Nick Wardle. This paper arose out of a workshop at the Banff Center for Arts and Creativity in Alberta, Canada, and we are very happy to thank the Banff International Research Station (BIRS) and its staff for their organisation and hospitality. We are grateful to Hans Dembinski for helpful comments.
ARB acknowledges that the publication was produced with funding from the Italian Ministry of University and Research under the Call for Proposals related to the scrolling of the final rankings of the PRIN 2022 call (Project title “Latent variable models and dimensionality reduction methods for complex data”, Project No. 20224CRB9E, CUP C53C24000730006, PI Prof. Paolo Giordani, Grant Assignment Decree No. 1401 adopted on 18.9.2024 by MUR).
IV acknowledges partial support by the US Department of Energy grant DE-SC0015592.
@@ A.1 The dimidiated Gaussian
Although the model is clearly unrealistic, it has some nice simple features. The quantile parameters provide the parameters of the pdf directly, as given by Equation (16)
@@
To determine the quantile parameters from the moments these equations have to be solved numerically. If we write $D={\sigma^{+}}-{\sigma^{-}}$ and $S={\sigma^{+}}^{2}+{\sigma^{-}}^{2}$, the equations
@@
${|\gamma|\over\sqrt{V^{3}}}={\pi+1\over\sqrt{(\pi-1)^{3}}}\approx 1.641,$ (20)
so asymmetries, in the $\gamma\over{V^{3/2}}$ sense, larger than this cannot be handled. But such cases are extreme.
@@ A.2 The distorted Gaussian
A more sophisticated procedure to handle the results of an OPAT analysis is to draw a parabola through the three points, as in Figure 20,
@@
The parameters $a$ and $b$ are given very simply from the quantile parameters: $a={1\over 2}({\sigma^{+}}+{\sigma^{-}})$ and $b={1\over 2}({\sigma^{+}}-{\sigma^{-}})$. The pdf is again the probability transform of the unit Gaussian, but using a single quadratic transform rather than two piecewise linear ones. Although this is more plausible than the two straight lines of the dimidiated Gaussian, care needs to be taken to include both arms of the parabola in evaluating the pdf. The support for $p(R)$ is limited by the minimum (or, for a negative skewness, the maximum) of the parabola. Where this occurs, there is a Jacobian spike, and this appears in Figure 20 around $R=22$.
If $\nu$ is distributed according to the standard Gaussian with $\mu=0$ and $\sigma=1$ and $x(\nu)=x_{0}+a\nu+b\nu^{2}$, then the pdf for $x$ is $p(x)={{\phi}(\nu)\over|x^{\prime}(\nu)|}={{\phi}({\nu})\over|a+2b\nu|}$, where $a$ and $b$ are as given in Equation (24), and $\phi(\nu)$ is the unit Gaussian. As this is a quadratic, there are two values of $\nu$ for a given $x$: $\nu={\sqrt{a^{2}+4bx}-a\over 2b}$ and ${-\sqrt{a^{2}+4bx}-a\over 2b}$. Except for small asymmetries and/or small deviations, where the contribution from the second can be neglected, the pdf for $x$ is the sum of the two.
@@ A.3 The “railway” Gaussian
The railway Gaussian density attempts to mitigate the drawbacks of the dimidiated Gaussian (the discontinuity in the middle) and of the distorted Gaussian (the Jacobian spike). This density is also based on a coordinate transformation. The curve drawn through the OPAT analysis points is a parabola in the [1,1][-1,1][ - 1 , 1 ] region which smoothly transitions into straight lines on both left and right sides of the region. These transition regions are highly reminiscent of the track transition curves in railroad tracks, and this similarity gives the density its name. The coordinate transformation is defined by the equation
@@
The R code (Appendix D) evaluates the moments of the railway Gaussian algebraically, whereas in the python/C++ code (Appendix C) the moments are evaluated numerically and, for numerical reasons, the values of $h$ used are restricted to the range [0.1,10.0][0.1,10.0][ 0.1 , 10.0 ].
@@ A.4 The double cubic Gaussian
Figure 22: The Double Cubic Gaussian. The panel on the left illustrates the coordinate transformation used to construct a double cubic Gaussian with parameters $M=0$, $\sigma^{+}=1$, $\sigma^{-}=0.6$. The panel on the right shows the corresponding density.
@@ A.5 The symmetric beta Gaussian
Figure 23: The Symmetric Beta Gaussian. The panel on the left illustrates the coordinate transformations used to construct densities shown on the right. In all cases, $M=0$, $\sigma^{+}=1$, $\sigma^{-}=0.6$.
@@ A.6 The quantile variable width Gaussian
An example QVW Gaussian density and the corresponding transformation of the standard normal are shown in Figure 24.
Figure 24: The QVW Gaussian. The panel on the right illustrates the density which corresponds to the asymmetric uncertainty result $0^{+1.0}_{-0.6}$. The panel on the left shows the corresponding coordinate transformation given by $x=q(\Phi(z))$.
@@ A.7 The Fechner distribution
Figure 25: The Fechner distribution. The panel on the right illustrates the probability density function which corresponds to the asymmetric uncertainty result $0^{+1.0}_{-0.7}$. (The result $0^{+1.0}_{-0.6}$ for which a number of other distributions are plotted in this appendix can not be represented by Fechner.) The panel on the left shows the corresponding coordinate transformation given by $x=q(\Phi(z))$.